15,000 is invested in a fund that has a return of 5%. how much money is invested in a fund with a 3% return if the toatl return on both investments is 1050

Answer by checkley79(3341)

.05*15000+.03X=1050
750+.03X=1050
.03X=1050-750
.03X=300
X=300/.03
X=10000 AMOUNT INVESTED @ 3%.
PROOF:
.05*15000+.03*10000=1050
750+300=1050
1050=1050